OpenStudy (anonymous):

the sum of 2 numbers is 9 and their difference is 6. What are the numbers?

Directrix (directrix):

Let x and y be the two numbers. Then, x + y = 9 and x - y = 6

Directrix (directrix):

x + y = 9 x - y = 6 --------- Add the above two equations. You can then solve for x and then y. Check to be sure the numbers do add to 9 and have a difference of 6. Post your answers here and someone will check, okay? @saber_doll

OpenStudy (anonymous):

@saber_doll if you follow @Directrix 's steps, it'll bring you to the right solution :) Do you know what the first step is?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

what is is?

OpenStudy (anonymous):

ok, you're going to solve for one equation in terms of the other. The two equations are: 1) x + y = 9 2) x - y = 6 so what you're going to do take the 1, ok? now, say that y=9-x (all i did was move x from one side to the other). Now you will substitute that value of y into the equation 2, so you will get x - (9-x) = 6. can you solve that?

OpenStudy (anonymous):

i'll try it now. thanks :)

OpenStudy (anonymous):

yep, let me know if you get it :)

OpenStudy (anonymous):

x+y+x-y=9+6 2x=15 x=7.5 7.5+y=9 y=9-7.5 y=1.5 I got it! is it right? :D @Khushal_Shah

OpenStudy (anonymous):

correct! nice job :) It's a little different than how I was saying, but it works out perfect in this situation :) Well done!

OpenStudy (anonymous):

Haha. Thanks alot you guys! :)))

OpenStudy (anonymous):

of course, anytime mate